OpenStudy (anonymous):

I need help with answer B and the other two could be checked thanks Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1620. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Donna spend on this trip?

OpenStudy (anonymous):

A: 50x+.6(200)=1620 B: C: 30 days

OpenStudy (anonymous):

@amistre64

OpenStudy (texaschic101):

A. correct.... B. 50x + .6(200) = 1620 --- distributive property 50x + 120 = 1620 -- subtract 120 from each side -- subtraction property ?? 50x = 1620 - 120 50x = 1500 -- divide by 50 -- division property x = 1500/50 x = 30 C. correct ..

OpenStudy (anonymous):

Thanks a lot these word problems are one after another and tend to be way too much after a while lol

OpenStudy (texaschic101):

I can understand that....after a while, they can get confusing...lol :)

OpenStudy (texaschic101):

thanks for the testimonial :)

OpenStudy (anonymous):

Thanks again ^~^ And no problem c:

OpenStudy (amistre64):

i would have prefered a more interactive solution process .... but yeah, word problems get to be a pain after awhile.

OpenStudy (amistre64):

also, the 'property' used is author specific in many cases.

OpenStudy (anonymous):

Yes a more interactive solution would have been nice but what I got was just as good. I didnt just look at the answer I tried doing it with the information as if I was trying to do it myself and checking afterward like everyone is meant to do if someone helps but they dont.